You wish to know the appropriate price for a security that matures in 6 years. The security has a payout of $24,777.46, which is paid in one lump sum at maturity. The interest rate associated with this class of security is 1.63%, and interest is compounded semiannually. What is the current price of this security?

Answer =

Cash Flow =

R =

N =
Answer #1

Current price of security is $ 22, 477.86

Current price of security is the present value of cash flows from security.
Present value of cash flow = Future value * Discount factor
= $ 24,777.46 * 0.90719
= $ 22,477.86
Working:
Discount factor = (1+i)^-n Where,
= (1+0.00815)^-12 i = 1.63%/2 = 0.00815
= 0.90718977 n = 6*2 = 12